OpenStudy (anonymous):

convert r^2=cos(theta) to rectangular coordinates

OpenStudy (anonymous):

I basicly multiply r on both sides. Which in turn it will be r*r^2=cos(theta) r In which cos(theta) r is "x" r*r^2=x And r^2 is x^2+y^2 And r is root(x^2+y^2) root(x^2+y^2)*x^2+y^2=x root(x^2+y^2)*x^2+y^2—x=0 I need to simplify? If i do can you help.

OpenStudy (anonymous):

I know the exponent for (x^2 +y^2) is 3/2 but why

OpenStudy (anonymous):

Nvm I found out why. Thanks useless people :D

OpenStudy (anonymous):

Just add exponents 1/2+1= 3/2